starting with the basic exponential function f(x)=3x (x is the power of 3), use transformation to sketch the graph of g(x)=3x-1 (x-1 is the power of 3)
Answer by lwsshak3(11628)

starting with the basic exponential function f(x)=3x (x is the power of 3), use transformation to sketch the graph of g(x)=3x-1 (x-1 is the power of 3)
**
If you graph the basic function, y=3^x, the curve will have a y-intercept of 1 (3^0=1)
For x>0, the curve will increase exponentially to ∞.
For x<0, the curve will approach zero with the x-axis being an asymptote to the curve.
..
For y=3^(x-1)
The entire curve is shifted one unit to the right.
For example, the y-intercept point of (0,1) has shifted to (1,1)
Otherwise, the shape of the curve remains the same with the x-axis still the asymptote for x<0